OpenStudy (anonymous):

Choose the correct slope of the line that passes through the points (0, 0) and (3, 30)

OpenStudy (anonymous):

each point is (x,y). so the first point is (x1,y1) and the second point is (x2,y2). Slope of a line is rise over run or rise/run. Rise is the change in the y direction and run is the change in the x direction. so slope = (y2-y1)/(x2-x1) = (30-0)/(3-0) = 30/3 = 10.